OpenStudy (anonymous):

lim q->1 logq^2(1+qw q-1/q+1)

OpenStudy (anonymous):

\[\lim_{q \rightarrow 1} \log _{q^2} (1+qw (\frac{ q-1 }{ q+1 }\]

OpenStudy (anonymous):

calculate with l'Hopital for a random \[w \in \mathbb{R} \]

hartnn (hartnn):

use the property of logarithm first \(\large \log_mn = \dfrac{\log n}{\log m}\) what u get ?

OpenStudy (anonymous):

\[\frac{ \log(1+qw \frac{ q-1 }{ q+1 }) }{ \log(q^2) }\]

OpenStudy (anonymous):

?

hartnn (hartnn):

yes, log q^2 = 2 log q , isn't it ? now you have the form 0/0 and can use L'hopitals, can you differentiate the numerator and denominator separately ?

OpenStudy (anonymous):

log q^2 = 2log q ....?? alright..

hartnn (hartnn):

yeah, \(\large \log x^m=m\log x\) thats a standard log property

OpenStudy (anonymous):

i get \[\frac{ 1 }{ 1+qw \frac{ q-1 }{ q+1 } } q^2\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

derivative of logx= 1/x right?

hartnn (hartnn):

do you know about chain rule ?

OpenStudy (anonymous):

oh right, lemme try this again

OpenStudy (anonymous):

so i just take the deriv. of log(..) times the deriv. (..) ?

OpenStudy (anonymous):

undefined, is that right?

OpenStudy (anonymous):

i get 0 for the num.

hartnn (hartnn):

hmm ? as you said, d/dx log x = 1/x so, in denominator, usingchain rule (1+qw(q-1/q+1)) times d/dq [(1+qw(q-1/q+1))]

hartnn (hartnn):

i don't get 0 for numerator...

OpenStudy (anonymous):

a little confused, i should review some things

OpenStudy (anonymous):

we started with \[\frac{ \log(1+qw \frac{ q-1 }{ q+1 }) }{ 2\log(q) }\]

OpenStudy (anonymous):

now i take the derivative of the numerator and denominator separately?

hartnn (hartnn):

yes,

hartnn (hartnn):

the numerator derivative will require some algebra

OpenStudy (anonymous):

i thought it was 0... i'll try again

OpenStudy (anonymous):

taking the derivative of the num. i get \[\frac{ 1 }{ 1+qw(\frac{ q-1 }{ q+1 }) } . (1+qw(\frac{ q-1 }{ q+1 })) \prime\]

OpenStudy (anonymous):

am i doing it right?

hartnn (hartnn):

yup, w is constant , it will be factored out of deivative. so, effectively, you have to apply quotient rule for \(\large w \dfrac{d}{dq}(\dfrac{q^2-q}{q+1})\)

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

applying the quotient rule i get \[\frac{ q+1-q-1 }{ (q+1)^2 }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i think i'm making a stupid mistake, just don't know what

hartnn (hartnn):

(2q-1) (q+1) - (q^2-q) = 2q^2 -q-1 -q^2+q = q^2 -1 is what i get for numerator

OpenStudy (anonymous):

\[(\frac{ f }{ g }) (a) = \frac{ f'(a)g(a)-f(a)g'(a) }{ (g(a))2 }\]

OpenStudy (anonymous):

using this i should get q^2-1?

OpenStudy (anonymous):

in the numerator

OpenStudy (anonymous):

i'll try again

hartnn (hartnn):

yes, derivative of numerator = 2q-1, derivative of numerator = 1

OpenStudy (anonymous):

derivative of num. is 2q-1 and 1? huh?

OpenStudy (anonymous):

oh, nevermind

hartnn (hartnn):

yeah, denominator, typo

OpenStudy (anonymous):

there's no way i'm getting 2q-1 in the num. ...i'm doing something very wrong with the algebra part

OpenStudy (anonymous):

and is it me, or is the algebra part difficult?

hartnn (hartnn):

num = q^2-q, denom = q+1 derivative of numerator times denominator - derivative of denominator times numerator = (2q-1) * (q+1) - 1 (q^2-q) = 2q*q +2q*1 -q*1 -1*1 -q^2 +q =2q^2-q^2 -q+q +2q -1 = q^2 +2q-1 sorry, i missed the 2q :(

OpenStudy (anonymous):

i'm so confused, you're showing how you found the num. using l'hopitals rule, applying quotient rule?

hartnn (hartnn):

wait i'll write all the steps.

OpenStudy (anonymous):

and q^2 +2q-1 is not equal to q^2-2q....

OpenStudy (anonymous):

that would help alot

hartnn (hartnn):

\(\large \frac{ 1 }{ 1+qw(\frac{ q-1 }{ q+1 }) } . (1+qw(\frac{ q-1 }{ q+1 })) \prime \\ \large = \frac{ 1 }{ 1+qw(\frac{ q-1 }{ q+1 }) } .w \dfrac{(2q-1)(q+1)-1(q^2-q)}{(q+1)^2}\\ \large = \frac{ 1 }{ 1+qw(\frac{ q-1 }{ q+1 }) } [w\dfrac{q^2+2q-1}{(q+1)^2}] \) so thats derivative of numerator, denominator is easy, log q^2 =2 log q its derivative = 2/q in denominator. so, finally we are here, \(\large = \frac{ q }{ 2(1+qw(\frac{ q-1 }{ q+1 })) } [w\dfrac{q^2+2q-1}{(q+1)^2}]\) see whether you have doubts ?

hartnn (hartnn):

now you can just plug in x=1

hartnn (hartnn):

***q=1

OpenStudy (anonymous):

yup, knew it. i made a stupid mistake

OpenStudy (anonymous):

is the answer 1 or w+1?

hartnn (hartnn):

\(\large = \frac{ q }{ 2(1+qw(\frac{ q-1 }{ q+1 })) } [w\dfrac{q^2+2q-1}{(q+1)^2}]\) with q=1 \(\large = \frac{ 1 }{ 2 } [w\dfrac{2}{4}]=\dfrac{w}{4}\) i get w/4

OpenStudy (anonymous):

alright, got it. thanks!

hartnn (hartnn):

welcome ^_^